Fing the circumference and area of each cirlce. Round to the nearest tenth

1.) The radius is 4cm
2.) The radius is 6cm.


Thanks so much

Answer by checkley71(8403)

CIR=2PIR OR C=2*3.14*4=25.12CM
AREA=PIR^2=3.14*4^2=50.24 CM^2
---------------------------------------------
CIR=2*3.14*6=37.68CM
AREA=3.14*6^2=113.04 CM^2